OpenStudy (anonymous):

What is the missing number in this sequence? 62 , 59 , 53 , ___ , 32

OpenStudy (anonymous):

A 47 B 44 C 42

OpenStudy (anonymous):

B) 44 62-59=3 59-53=6 53-x=9 x-32=12